OpenStudy (anonymous):

Translate to a system of equations and solve Molly is 4 years younger than Heidi. In 3 years, Heidi will be twice as old as Molly. How old are they now ?

OpenStudy (anonymous):

Molly is 4 years younger than Heidi: M = H + 4 In 3 years, Heidi will be twice as old as Molly: H + 3 = 2(M + 3) Now, solve...

OpenStudy (anonymous):

SOrry...M = H - 4